# Khmer Semantic Search Engine: Digital Information Access and Document Retrieval

### v1

@@ Abstract
The search engine process is crucial for document content retrieval. For Khmer documents, a tool is needed to extract essential keywords. Despite the daily generation of significant Khmer content, Cambodians struggle to find necessary documents due to the lack of an effective semantic searching tool. Even Google does not deliver high accuracy for Khmer content. Semantic search engines improve search results by employing advanced algorithms to understand various content types. With the rise in Khmer digital content—such as reports, articles, and social media feedback—enhanced search capabilities are essential. This research proposes the first Khmer Semantic Search Engine (KSE), designed to improve traditional Khmer search methods. Utilizing semantic matching techniques and formally annotated semantic content, our tool extracts meaningful keywords from user queries performs precise matching, and provides the best matching offline documents and online URL documents. We propose two semantic search frameworks based on keyword extraction and semantic search matching. Additionally, we developed tools for data preparation, including document addition and manual keyword extraction. To evaluate performance, we created a ground truth dataset and discussed issues related to searching and semantic search. Our findings show how understanding search term semantics can lead to more accurate results.
@@ I Introduction
This paper presents the research design, implementation, and evaluation of KSE. It proposes two semantic search systems based on keyword extraction and semantic search matching. (See detail in Fig. 1)
Additionally, it describes the development of tools for data preparation, including document addition and manual keyword extraction. The evaluation of KSE’s performance, using a ground truth dataset, demonstrates how an understanding of the semantics of search terms can significantly improve search accuracy.
@@
3. Comprehensive Data Preparation Tools: We developed tools for data preparation, including automatic and manual keyword extraction, which ensure high-quality input data for the search engine.
@@
In the following sections, we will dive into the introduction, objectives, and problems associated with building an effective Khmer semantic search engine. We will also review related literature, propose the KSE architecture, and discuss the implementation and evaluation methodologies. Finally, we will conclude with the findings and future perspectives of this research.
@@ IV-A Models of Offline and Online Processing
IV-B Input Modules
1) Query Input: Users enter their search queries in Khmer through a user-friendly graphical user interface (GUI). This input is crucial for initiating the search process. The GUI is designed to be intuitive and accessible, encouraging users to engage with the search engine without any technical difficulties. For example, a user might enter the query "Cambodia".
2) Document Input: Documents, including reports, articles, and social media feedback, are added to the system. These documents are processed and indexed for efficient retrieval. The input can come from various sources, such as uploading files, entering text directly, or scraping web content. This ensures comprehensive coverage of content.
IV-C Output Modules
1) Search Results: The system outputs the most relevant search results based on the user’s query. These results include URLs and a relevance score, providing users with a clear indication of the usefulness of each result. For instance, if the query is "Cambodia" the results might include links to government reports, historical articles, and relevant social media posts.
2) Result Ranking: The search results are ranked based on their relevance to the query. The relevance scores are calculated by considering the semantic similarity between the query and document keywords. The ranked search results are then displayed to the user through the GUI. The most relevant results appear at the top, helping users quickly find the information.
IV-D Detailed Processes
@@
IV-D1 Process 1: Index the Editable Document
Users input documents into the system for indexing purposes. These indexing processes accelerate the document searching process when users search for relevant documents. This process contains two steps:
Step 1: Keyword Extraction & Indexation: The editable document is used as input into the “Keyword Extraction & Indexation” module. This module extracts a few main keywords from the document and indexes the input documents with those extracted keywords. The main keywords significantly explain the content of the document. For example, if a document is about Cambodian history, keywords might include "Cambodia," "history," and "Khmer Empire."
Step 2: Storage: The extracted keywords from Step 1 and the input document are then stored in the database. This indexed information allows for fast retrieval when a search query is made.
IV-D2 Process 2: Semantic Searching Process
This process is responsible for finding the most suitable document by comparing it to the user’s input sentence. This process contains six steps:
Step 1: Query Input: The user inputs a search sentence into the system through a GUI. For example, "What is the history of Angkor Wat?"
Step 2: Keyword Extraction & Indexation: This module performs keyword extraction from the query sentence. In the case of documents, it stores the document and indexes it in the database.
Step 3: Semantic Matching: This module takes the keywords extracted from the user’s query and the keywords extracted from the document, then performs semantic matching with the support of domain ontology and a dictionary. It outputs the document most relevant to the user’s query. For example, it might match the query "Angkor Wat" with documents containing detailed historical information about the temple.
Step 4: Result Ranking: After receiving the matching result from Step 3, this module ranks the results based on similarity scores. The most relevant one appears at the top, followed by less similar results.
Step 5: Result Display on the GUI: The results generated from Step 4 are displayed to users through the GUI interface. Users see a list of documents with relevance scores and links to the full text.
Step 6: Explore the Result: Through the GUI provided by the system, users can explore the details of the content of the result document or link. For instance, they might click on a link to read a full report on the history of Angkor Wat.
@@ V-A Stop word dictionary
By removing these stop words from the article or document, we can extract the more meaningful keywords and concepts. These extracted keywords then serve as the basis for the semantic search, allowing the system to understand the core meaning and context of the content and return more relevant and precise search results [24].
@@ V-E Evaluation methods
We evaluated our Khmer Semantic Search Engine (KSE) using two main criteria: keyword extraction and search result ranking.
@@ V-E1 Keyword Extraction
V-E2 Score ranking
@@ VI-A Comparison of Keyword Extraction Methods
TABLE IV: Comparison of Search Results for Top 5 Documents.
@@
TABLE V: Results of Our Proposed KSE Across Documents.
@@
VI-B Comparison of Search Results Ranking
Table IV presents the input keywords generated by the tool and the top 5 documents returned for different search queries. This table provides a clear view of how the search engine prioritizes documents based on the relevance of the extracted keywords. By comparing these results with the manually curated top 5 documents, we can determine the accuracy and effectiveness of the search engine.
Additionally, Table V provides the calculated scores of the search testing rankings, including precision, recall, and F1-score, compared to the manually curated documents. Precision measures the proportion of relevant documents retrieved out of the total documents retrieved by the search engine. Recall evaluates the proportion of relevant documents retrieved out of the total relevant documents available. The F1-score, being the harmonic mean of precision and recall, offers a balanced measure of the search engine’s performance.
The results of our testing, as summarized in Table V, demonstrate that our search engine achieves an average F1-score of 0.75, indicating a high level of effectiveness in providing accurate and relevant search results for Khmer language content. This performance is above average and showcases the potential of our Khmer Semantic Search Engine (KSE) as a significant advancement over traditional search engines for the Khmer language.

### v2

# Khmer Semantic Search Engine (KSE): Digital Information Access and Document Retrieval

@@ Abstract
The search engine process is crucial for document content retrieval. For Khmer documents, an effective tool is needed to extract essential keywords and facilitate accurate searches. Despite the daily generation of significant Khmer content, Cambodians struggle to find necessary documents due to the lack of an effective semantic searching tool. Even Google does not deliver high accuracy for Khmer content. Semantic search engines improve search results by employing advanced algorithms to understand various content types. With the rise in Khmer digital content—such as reports, articles, and social media feedback—enhanced search capabilities are essential. This research proposes the first Khmer Semantic Search Engine (KSE), designed to enhance traditional Khmer search methods. Utilizing semantic matching techniques and formally annotated semantic content, our tool extracts meaningful keywords from user queries, performs precise matching, and provides the best matching offline documents and online URLs. We propose three semantic search frameworks: semantic search based on a keyword dictionary, semantic search based on ontology, and semantic search based on ranking. Additionally, we developed tools for data preparation, including document addition and manual keyword extraction. To evaluate performance, we created a ground truth dataset and addressed issues related to searching and semantic search. Our findings demonstrate that understanding search term semantics can lead to significantly more accurate results.
@@ I Introduction
This paper presents the research design, implementation, and evaluation of the Khmer Semantic Search Engine (KSE). It proposes three semantic search frameworks: semantic search based on a keyword dictionary, semantic search based on ontology, and semantic search based on ranking.
Additionally, it describes the development of tools for data preparation, including document addition and manual keyword extraction. The evaluation of KSE’s performance, using a ground truth dataset, demonstrates how understanding the semantics of search terms can significantly improve search accuracy.
@@
3. Comprehensive Data Preparation Tools: We developed tools for data preparation, including automatic and manual keyword extraction, and ontology-based semantic enrichment to ensure high-quality input data for the search engine.
@@
In the following sections, we will delve into the introduction, objectives, and challenges associated with building an effective Khmer semantic search engine. We will also review related literature, propose the KSE architecture, and discuss the implementation and evaluation methodologies. Finally, we will conclude with the findings and future perspectives of this research.
@@ IV-A Models of Offline and Online Processing
IV-B Detailed Processes
@@
1) Input Modules
Query Input: Users enter their search queries in Khmer through a user-friendly graphical user interface (GUI). This input is crucial for initiating the search process. The GUI is designed to be intuitive and accessible, encouraging users to engage with the search engine without any technical difficulties. For example, a user might enter the query "Cambodia".
Document Input: Documents, including reports, articles, and social media feedback, are added to the system. These documents are processed and indexed for efficient retrieval. The input can come from various sources, such as uploading files, entering text directly, or scraping web content. This ensures comprehensive coverage of content.
2) Output Modules
Search Results: The system outputs the most relevant search results based on the user’s query. These results include URLs and a relevance score, providing users with a clear indication of the usefulness of each result. For instance, if the query is "Cambodia" the results might include links to government reports, historical articles, and relevant social media posts.
Result Ranking: The search results are ranked based on their relevance to the query. The relevance scores are calculated by considering the semantic similarity between the query and document keywords. The ranked search results are then displayed to the user through the GUI. The most relevant results appear at the top, helping users quickly find the information.
IV-C Semantic Search Matching Process
The semantic search matching process in our Khmer Semantic Search Engine (KSE) involves three main components: semantic search based on keywords, semantic search based on tourism ontology for Cambodia, and semantic search based on ranking. Each component is designed to enhance the accuracy and relevance of the search results through detailed and technically robust processes.
1)Semantic Search Based on Keywords
The first component involves a sophisticated keyword extraction process from user queries. This begins with preprocessing steps such as:
Tokenization: The query is split into individual words or tokens. For example, the query "best cultural sites in Phnom Penh" is tokenized into ["best", "cultural", "sites", "in", "Phnom", "Penh"].
Stop-Word Removal: Common Khmer stop words are removed to focus on significant terms. This step uses a predefined list of stop words specific to the Khmer language, as detailed in [8]. For instance, Khmer stop words like (in), (of), (has), and (is) are filtered out. Example: the query (waterfalls in Phnom Penh) would be reduced to [(waterfalls), (Phnom Penh)].
Stemming and Lemmatization: Words are reduced to their base or root form. For instance, "sites" becomes "site". This step ensures that variations of a word are considered equivalent.
TF-IDF Calculation: Term Frequency-Inverse Document Frequency (TF-IDF) is used to weigh the importance of each term within the query relative to its frequency in the document corpus.
Keyword Matching: The extracted and weighted keywords are then used to search the indexed documents. Each document in the database has been pre-processed in a similar manner, with keywords extracted and stored in an index for fast retrieval. The system compares the keywords from the user query with those in the document index to find potential matches.
2) Semantic Search Based on Tourism Ontology for Cambodia
The second component leverages a domain-specific ontology to enhance the semantic understanding of user queries. The tourism ontology for Cambodia includes detailed information about various tourist destinations, cultural landmarks, historical sites, and other points of interest in Cambodia. The ontology provides a structured representation of domain knowledge, including the relationships between different entities.
The ontology is developed using the Web Ontology Language (OWL) and consists of multiple classes and properties that capture the complexities of Cambodian tourism. Specifically, the ontology includes:
Entities: These include tourist attractions (e.g., Angkor Wat, Royal Palace), types of attractions (e.g., temples, museums), and locations (e.g., Phnom Penh, Siem Reap). Each entity is described with properties such as historical significance, visitor reviews, operational hours, and geographical coordinates.
Relationships: Relationships define how entities interact. For example, "Angkor Wat" (entity) is a "temple" (type) located in "Siem Reap" (location). The ontology also includes hierarchical relationships, such as "Siem Reap" is a part of "Cambodia".
Ontology-Based Query Expansion: When a user enters a query, the system uses the ontology to expand the query semantically. For instance, a query for "temples in Phnom Penh" might expand to include synonyms and related terms like "wats" and specific temple names in Phnom Penh. The system leverages synonyms and related entities defined in the ontology to ensure comprehensive query coverage.
Ontology Details: The ontology comprises over 500 entities, covering major and minor tourist attractions, cultural sites, and events. It includes approximately 1000 relationships that map these entities in various contexts, providing a rich semantic network for query expansion. For example, a query for "temples in Phnom Penh" might be expanded to include "Wat Phnom" and "Wat Botum", both specific temple names in Phnom Penh.
Semantic Similarity Calculation: The system calculates the semantic similarity between the expanded query and the documents using ontology-based metrics. This might involve techniques such as cosine similarity on the vector representations of entities and their relationships, ensuring that documents related to the context of the query are prioritized.
3) Semantic Search Based on Ranking
The final component focuses on ranking the retrieved documents to present the most relevant results to the user. This process combines both keyword-based and ontology-based relevance and involves several steps:
Relevance Scoring: Documents are scored based on both keyword relevance and semantic relevance. Keyword relevance is determined by the presence and frequency of query keywords in the document, as calculated using TF-IDF scores. Semantic relevance is assessed by comparing the ontology-based context of the query with the document content. This involves measuring the semantic similarity between the entities and relationships identified in the query and those in the documents.
Weighted Scoring Algorithm: The overall score for each document is a weighted combination of keyword relevance and semantic relevance. This approach ensures that both the direct relevance of keywords and the deeper semantic connections are considered, providing a balanced and accurate ranking of documents.
Popularity Metrics: In addition to relevance scores, documents are also ranked based on popularity metrics such as click-through rates, user ratings, and the number of views. These metrics provide insight into how users interact with the documents, indicating their perceived value and relevance. To incorporate these metrics into the final ranking score, we normalize the data to ensure comparability. For example, click-through rates might be scaled to a range between 0 and 1, user ratings could be averaged, and view counts normalized based on the highest and lowest values in the dataset. By including popularity metrics, we ensure that frequently accessed and highly rated documents receive an appropriate boost in their ranking, reflecting their utility and relevance to users.
Ranking Adjustments: The system applies final adjustments to the ranking based on additional factors such as the recency of the document, user-specific preferences, and contextual relevance. For instance, more recent documents or those with higher user engagement might receive a slight boost in their ranking. This ensures that the most relevant and timely information is presented to the user.
Presentation of Results: The top-ranked documents are then presented to the user in a list, with the highest-scoring documents appearing first. Each result includes a brief snippet highlighting the relevant keywords and context to help the user quickly identify the most pertinent information. This presentation aims to enhance the user experience by making it easier to find the desired information quickly.
By integrating these three components, the KSE provides a comprehensive and accurate search experience tailored to the specific needs of Khmer language users and the domain of Cambodian tourism. This multi-layered approach ensures that the search engine can handle complex queries and deliver relevant, high-quality results effectively.
@@ V-A Stop word dictionary
By removing these stop words from the article or document, we can extract the more meaningful keywords and concepts. These extracted keywords then serve as the basis for the semantic search, allowing the system to understand the core meaning and context of the content and return more relevant and precise search results.
@@ V-E Evaluation Methods
We evaluated our Khmer Semantic Search Engine (KSE) using three main criteria: keyword extraction, ontology-based semantic search, and search result ranking.
@@ V-E1 Keyword Extraction
V-E2 Ontology-Based Semantic Search
To evaluate the effectiveness of the ontology-based semantic search, we used similar metrics: precision, recall, and F1 score. The ontology-based approach leverages domain knowledge to expand and refine queries, aiming to retrieve more contextually relevant documents. The evaluation process involved:
  • Precision: The proportion of documents retrieved that are relevant based on the expanded query.
  • Recall: The proportion of relevant documents retrieved out of the total relevant documents available.
  • F1 Score: The harmonic mean of precision and recall.
Precision=Number of relevant documents retrievedTotal number of documents $\text{Precision}=\frac{\text{Number of relevant documents retrieved}}{\text{Total number of documents retrieved}}$
Recall=Number of relevant documents retrievedTotal number of relevant $\text{Recall}=\frac{\text{Number of relevant documents retrieved}}{\text{Totalnumber of relevant documents}}$
[…]
These metrics help us assess how well the ontology-based search improves the relevance of search results by understanding the semantic context of queries.
V-E3 Score Ranking
@@
These combined evaluation methods allow us to comprehensively assess the performance of the KSE in extracting keywords, understanding semantic context through the ontology, and ranking search results effectively.
@@ VI-A Comparison of Keyword Extraction Methods
VI-B Comparison of Ontology-Based Semantic Search
To evaluate the performance of the ontology-based semantic search, we used a similar approach as with keyword extraction. The ontology provides a structured representation of domain knowledge, allowing the search engine to understand the context of queries better. This section compares the search results obtained using ontology-based methods with those from manual extraction.
The ontology-based semantic search significantly improves the search accuracy by leveraging the structured domain knowledge provided by the ontology. The expansion of user queries using ontology allows the search engine to understand the context and retrieve documents that are more relevant, even if they do not contain the exact search terms. For example, a search for "temples in Phnom Penh" not only retrieves documents that mention "temples" and "Phnom Penh" but also includes documents that discuss "wats" and specific temples like "Wat Phnom" and "Wat Botum."
This approach addresses the limitations of keyword-based search methods, which often miss contextually relevant documents. By incorporating semantic understanding, the ontology-based search can handle complex queries and provide more accurate and comprehensive search results.
Additionally, the use of relationships within the ontology helps in disambiguating terms that might have multiple meanings. For instance, "Angkor" could refer to various things, but within the tourism context, it most likely refers to the Angkor Wat temple complex. By using the ontology, the system can infer the correct meaning based on the context provided by the user query.
Table IV shows how the ontology-based approach enhances the relevance of search results by including contextually related terms and entities. The evaluation metrics include precision, recall, and F1-score, similar to the keyword extraction evaluation.
Overall, the ontology-based semantic search demonstrates its effectiveness in enhancing the relevance and accuracy of search results for Khmer language content. This approach not only improves the search experience for users but also sets a foundation for further advancements in semantic search technologies for low-resource languages like Khmer.
TABLE IV: Comparison of Ontology-Based Semantic Search
Document ID Manual Keywords Ontology-Based Keywords Precision Recall F1
1 kw1, kw2, kw3 kw1, kw2, kw4 0.90 0.85 0.87
2 kw4, kw5, kw6 kw4, kw5, kw7 0.85 0.80 0.82
3 kw7, kw8, kw9 kw7, kw8, kw10 0.88 0.84 0.86
4 kw10, kw11, kw12 kw10, kw11, kw13 0.92 0.87 0.89
5 kw13, kw14, kw15 kw13, kw14, kw16 0.89 0.86 0.87
VI-C Comparison of Search Results Ranking
Table V presents the input keywords generated by the tool and the top 5 documents returned for different search queries. This table provides a clear view of how the search engine prioritizes documents based on the relevance of the extracted keywords. By comparing these results with the manually curated top 5 documents, we can determine the accuracy and effectiveness of the search engine.
Additionally, Table VI provides the calculated scores of the search testing rankings, including precision, recall, and F1-score, compared to the manually curated documents. Precision measures the proportion of relevant documents retrieved out of the total documents retrieved by the search engine. Recall evaluates the proportion of relevant documents retrieved out of the total relevant documents available. The F1-score, being the harmonic mean of precision and recall, offers a balanced measure of the search engine’s performance.
The results of our testing, as summarized in Table VI, demonstrate that our search engine achieves an average F1-score of 0.75, indicating a high level of effectiveness in providing accurate and relevant search results for Khmer language content. This performance is above average and showcases the potential of our Khmer Semantic Search Engine (KSE) as a significant advancement over traditional search engines for the Khmer language.
TABLE V: Comparison of Search Results for Top 5 Documents
@@ VI-C Comparison of Search Results Ranking
TABLE VI: Results of Our Proposed KSE Across Documents
@@ VII Conclusion
Overall, KSE’s development marks a pivotal advancement in enhancing information retrieval for the Khmer language. By overcoming language-specific challenges and incorporating advanced semantic matching techniques, our system provides a robust foundation for further innovation. Future work will focus on integrating more sophisticated machine learning models and expanding the ontology to cover broader domains, ultimately aiming to provide even more accurate and contextually relevant search results for Khmer-speaking users.